OpenStudy (anonymous):

Which of the following statements best describes the effect of replacing the graph of f(x) with the graph of f(x) − 3? The graph shifts 3 units up. The graph shifts 3 units down. The graph shifts 3 units left. The graph shifts 3 units right. i will give a medal

OpenStudy (anonymous):

please help i will give a medal out and fan

OpenStudy (mathmale):

Hello. Please see page 12 (Transformations) of the following: http://www2.sunysuffolk.edu/pestiej/MAT124_transformations.pdf I'd be happy to answer questions about these Transformations after you've taken a look at this material.

OpenStudy (anonymous):

@mathmale i went to page 12 i need help and i dont get this at all

OpenStudy (mathmale):

There are four basic transformations: Starting with a basic graph, we can shift (move) the graph up or down, to the left or to the right. Up or down involves the y-axis; left or right involves the y axis. Let's build upon what you already know about transformations. What is that?

OpenStudy (anonymous):

idk anything about transformation

OpenStudy (mathmale):

You are given the function y = f(x) − 3, as well as the graph of y = f(x). Referring to the summary on page 12, which of the four cases on page 12 is closest to the problem you're working on now?

OpenStudy (mathmale):

My job is to help you solve the given problem yourself, not to do it for you. Please answer my question (above).

OpenStudy (anonymous):

ok let me look

OpenStudy (anonymous):

Reflection: y = −f(x) Reflect the graph of y = f(x) in the x-axis is it this?

OpenStudy (anonymous):

@mathmale am i right?

OpenStudy (mathmale):

You are given y = f(x) - 3. Compare that to the four cases that follow: y=f(x) + k y=f(x) - k y=f(x-a) y=f(x+a) Which of these four is closest to the problem given you?

OpenStudy (anonymous):

y=f(x) - k?

OpenStudy (mathmale):

Yes. What does the summary on page 12 say happens when you start with the graph of f(x) and graph the equation y=f(x)-3?

OpenStudy (mathmale):

@bashbrothers : Please respond.

OpenStudy (mathmale):

I see you are new to OpenStudy. Please, as a matter of courtesy to those trying to help you, either stick with your conversation with that person or let him/her know if you are no longer interested in continuing the discussion.